What is the pH of a 1.00 M solution of disodium sulfate (Na2SO4)?
You can calculate the Kb from the Ka values given in the question 1. (Ka1 = large
Ka2 = 1.2 x 10-2)

  1. 7.94
  2. 8.02
  3. 6.01
  4. 7.96
Answer #1

Na2SO4 ( aq ) \rightarrow 2 Na+ ( aq ) + SO42- ( aq )

SO42- ( aq ) + H2O ( l ) \rightleftharpoons HSO4- ( aq ) + OH- ( aq )

Kb ( SO42- ) = Kw / Ka2 = 10–14 / 1.2 x 10-2 = 8.33 x 10-13

calculations So, (aq) + H₂O = HSO i to the tre (1-x) ta tre -13 Kp = 8.33 810 = 2 x I- 1 - 2 = [01] =9.12 68 x 10 mg = -109 (9.1268 X10 )